OpenStudy (anonymous):

Solve the differential: y''+[(1/x)*(y')]-[(1/x^2)*y]=ln(x) What would be the best way to approach a problem like this? Variation of parameters?

OpenStudy (amistre64):

theres a long way and an even longer way

OpenStudy (amistre64):

a wronskian appraoch is prolly the simplest and shortest longer way

OpenStudy (amistre64):

whats our homogenoues solution using the characteristic eq?

OpenStudy (anonymous):

can you explain what you mean by using the wronskian? just to find the independent solutions?

OpenStudy (amistre64):

i can, but its really the result of using variation of parameters, the longer method

OpenStudy (amistre64):

first find the homog part assuming x^r is a solution

OpenStudy (amistre64):

y''+[(1/x)*(y')]-[(1/x^2)*y]=ln(x) rewrite it as; x^2 y''+x y'- y = x^2 ln(x)

OpenStudy (amistre64):

assuming x^m is a solution to the couchy

OpenStudy (anonymous):

ok im following

OpenStudy (amistre64):

y = x^m y' = x x^(m-1) y'' = x(x-1) x^(m-2) sub those in to find the homog solution, and our y1 y2 generals that we will need in the wronskian

OpenStudy (amistre64):

then we set up the Wronskian, or rather i set it up, like this and take a pseudo cross \[\begin{vmatrix}W_1&W&W_2\\y_1&0&y_2\\y'_1&ln(x)&y'_2\end{vmatrix}\]

OpenStudy (amistre64):

\[W_1=-y_2ln(x)\]\[W=y_1y'_2-y_2y'_1\]\[W_2=y_1ln(x)\] our particular solution then turns out to be \[yp_1 = y_1\int\frac{W_1}{W}\]\[yp_2 = y_2\int\frac{W_2}{W}\]

OpenStudy (anonymous):

Can you explain why you set it up that way? isn't a Wronskian supposed to look like this? \[\left[\begin{matrix}y1 & y2 & y3 \\ y1' & y2' & y3' \\ y1(n-1) & ... & y3(n-1) \end{matrix}\right]\]

OpenStudy (amistre64):

my set up is only a means to organize the information

OpenStudy (amistre64):

and no, your example is not the Wronskian method

OpenStudy (amistre64):

similar yes, but not organized correctly

OpenStudy (anonymous):

I didn't learn that way yet :\

OpenStudy (amistre64):

the Wronskian is the cramer rule result of doing it the longer way

OpenStudy (anonymous):

I see, I will have to read up on those two methods since we haven't covered them yet. Do you have any good resources or suggestions for reading?

OpenStudy (amistre64):

i havent seen any one set it up like i do, simply because i develop my own style. but this usually is coherent enough to follow http://tutorial.math.lamar.edu/Classes/DE/Wronskian.aspx

OpenStudy (anonymous):

Ok thank you! I will try to follow that approach and see if I can translate it to the problem.. But you say this is possible with variation of parameters as well?

OpenStudy (amistre64):

yes, since this is just a restructring of the variation of parameters

OpenStudy (anonymous):

one more question - why did you multiply by x^2 in the beginning?

OpenStudy (amistre64):

in order to solve the homogenous part; we dont need the fractions

OpenStudy (anonymous):

The Wronskian is related!!! I see, we didnt learn that yet.. hum

OpenStudy (amistre64):

we can go thru the longer version just as well

OpenStudy (anonymous):

Wait I'm still trying to understand everything you did, since I've seen a similar form used briefly by my teacher but not well explained - are you assuming a power law when you write: y = x^m; y' = (x)(x^m-1); y'' = (x-1)(x^m-2)

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

And then you plug into the original differential to use the W approach?

OpenStudy (amistre64):

the functions of x as coeefs tells me this is a couchy form of a diffyQ

OpenStudy (amistre64):

\[x^2(m(m-1))x^{m-2}\]\[+x(m-1)x^m\]\[-x^m\] \[x^m(m^2-m+m-1-1)=0;\ m^2=0\] we have a double root it looks like so:\[y=c_1x^0+c_2x^0ln(x)+yp\]

OpenStudy (anonymous):

Ok - just read the wiki and I follow your approach - but how to solve for yp then?

OpenStudy (anonymous):

Using this approach?

OpenStudy (amistre64):

assume c1 and c2 are functions of x; A and B \[yp=A+Bln(x)\] \[yp'=A'+B'ln(x)+\frac{B}{x};\ A'+B'ln(x)=0\]\[yp'=\frac{B}{x}\] \[yp''=\frac{B'}{x}+\frac{xB'-B}{x^2} \] with any luck, when we use these in the original, things get simplified :)

OpenStudy (amistre64):

\[x^2(\frac{B'}{x}+\frac{xB'-B}{x^2})+x(\frac{B}{x})-(A+Bln(x))]=x^2 ln(x)\] \[B'x+xB'-B+B-A-Bln(x) =x^2 ln(x)\] \[2B'x-A-Bln(x) =x^2 ln(x)\] seems im rusty on the longest of the longer version lol

OpenStudy (anonymous):

\[y _{h}=[x ^{2}(m-1)(x^{m-2})]+[x(m*x^{m-1})]-x^{m}=0\]

OpenStudy (anonymous):

thats what I get for the homogenous solution - which is different than your solution -_-

OpenStudy (amistre64):

you didnt derive x^m correctly for each spot

OpenStudy (anonymous):

x^m; m*x^m-1; and (m-1)(x^m-2)

OpenStudy (anonymous):

for y y' and y''

OpenStudy (amistre64):

\[y=x^m\]\[y'=m\ x^{(m-1)}\]\[y''=m(m-1)\ x^{(m-2)}\]

OpenStudy (anonymous):

ohhh i forgot an m

OpenStudy (amistre64):

we are just using the power rule

OpenStudy (anonymous):

Ok so then you factor x out and make an expression for m - then solve for the roots of m to get your homogenous solution right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

how do you group the terms that have x to some m power? - i dont think i've factored an expression like this before

OpenStudy (anonymous):

I get x^m+1 on the outside - but idk if im diong it right

OpenStudy (amistre64):

\[x^2 * x^{m-2}=x^{m-2+2}\to\ x^m\]

OpenStudy (amistre64):

all the xpart go away and we are left with an x^m to factor out that never goes zero so we can ignore it

OpenStudy (amistre64):

m(m-1) + m - 1 = 0

OpenStudy (amistre64):

m^2 - 1 = 0 when m=1 or -1

OpenStudy (amistre64):

looks like a mismathed at the start of this post also :)

OpenStudy (anonymous):

yes i thought it was just a typo but i guess it carried to the solution

OpenStudy (anonymous):

this is what i have so far worked out

OpenStudy (anonymous):

\[y_{h}=x^{m}(m(m-1))+x^{m-1}(m)-x^{m}\]

OpenStudy (anonymous):

(btw do you know if this board uses latex for equations?)

OpenStudy (amistre64):

it codes latex

OpenStudy (anonymous):

ohh thats awesome - bc ive been meaning to learn, now I will have to use this all the time (my first time on this board)

OpenStudy (amistre64):

your middle should be x^m as well; not x^(m-1)

OpenStudy (anonymous):

\[x(y')=x(m*x^{m-1}) since y=x^{m} and y'=m*x^{m-1}]

OpenStudy (amistre64):

x*x^(m-1) = x^m

OpenStudy (anonymous):

er\[(x(y')=x(m*x^{m-1}) since y=x^{m} and y'=m*x^{m-1}\]

OpenStudy (amistre64):

\text{} writes in latex

OpenStudy (anonymous):

ah ok noted

OpenStudy (amistre64):

\[this and that\text{ : this and that}\]

OpenStudy (anonymous):

\[x(y′)=x(m∗xm−1)\text{ since, }y=xm\text{ and, }y′=m∗xm−1\]

OpenStudy (anonymous):

forgot to put in exp when copy paste

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

I got x^m now though

OpenStudy (anonymous):

x*x^{m-1} = x^{m} right?

OpenStudy (amistre64):

yes, simplifying to m x^m

OpenStudy (anonymous):

Yup - ok so the general solution to this then is y = Cx^m

OpenStudy (amistre64):

yes, but:\[y=c_1x^{-1}+c_2x^1+yp\]

OpenStudy (anonymous):

so now i can use variation of parameters or solve for the coefficients

OpenStudy (amistre64):

yes yp = Ax^-1 + Bx yp' = A'x^-1 -Ax^-2 +B'x + Bx' ; the A' B' stuff sums to zero yp' = -Ax^-2 + B yp'' = -A'x^-2 +2Ax^-3 + B'

OpenStudy (anonymous):

yup yup i see

OpenStudy (amistre64):

x^2 yp'' = -A' +2A/x + B'x^2 +x yp' = -A/x + Bx -yp = -A/x - Bx -------------------------- x^2 ln(x) = -A' + B'x^2

OpenStudy (amistre64):

now we setup a system of eqs; we assumed; A'/x + B' x = 0 and now; A'/x + B' x = x^2 ln(x)

OpenStudy (amistre64):

solve for A' B' and then integrate them up tp A and B

OpenStudy (anonymous):

how does A'x^-1 go to zero in : yp' = A'x^-1 -Ax^-2 +B'x + Bx' ; the A' B' stuff sums to zero

OpenStudy (anonymous):

A'x^-1 + B'x = 0 ... yes but can you add them?

OpenStudy (amistre64):

since c1 x^-1 + c2x = 0; its really just the result of the yh part since c1=A and c2 = B

OpenStudy (anonymous):

so yh'=0 then?

OpenStudy (amistre64):

yes, thats the definition of homogenous equations; this = 0

OpenStudy (amistre64):

and since c1 is arbitrary and c2 as well; A' and B' makes no difference; it just goes to zero

OpenStudy (anonymous):

ah subtle - but straightforward, ok i believe you

OpenStudy (anonymous):

ok still working through it, trying to solve for coeffs

OpenStudy (amistre64):

as a result, we can cancel them from the yp part to determine the particular results

OpenStudy (anonymous):

Ok, so I got this: x^2 ln(x) = -A' + B'x^2 --> but you said that: we assumed; A'/x + B' x = 0 and now; A'/x + B' x = x^2 ln(x)

OpenStudy (amistre64):

one is a result of the yh part ; which equals 0 the other is a result of the yp part; which equals x^2 ln(x)

OpenStudy (anonymous):

So i use these two equations to solve for the coefficients? What about the x^2*ln(x) = -A' + B'*x^2

OpenStudy (anonymous):

what was the point of getting tha result?

OpenStudy (amistre64):

in order to put it all together since the general solution is the sum of the yh and the yp parts

OpenStudy (amistre64):

but we have to determine A and B for the general solution, not A' and B'; so we have to solve this system for A' and B' in order to integrate A' to A, and B' to B

OpenStudy (anonymous):

Yes i understand that part but now I am confused with which equations to use to solve for A' and B', it is not A'x+B'x = 0 -A'+B'x^2 = ln(x)*x^2

OpenStudy (amistre64):

i was just checking the wolf to make sure we was on the right track

OpenStudy (anonymous):

Ok - i hope it is right, bc ti makes sense to me ha